OpenStudy (anonymous):

The third term of an arithmetic sequence is 3 and the seventh term is 15. What is the 17th term?

OpenStudy (anonymous):

ok so what is added to each term to make the next term? take the distance of the terms 7 - 3 = 4 take the difference of the terms 15 - 3 = 12 difference / distance = 12/4 = 3 each term is the previous term + 3 I think you can get it from there

OpenStudy (anonymous):

so for example term 3 is 3 and term 4 is 3 + 3 which is 3*2 you can write that as 3 * (number of term you are on minus 2), often written as 3 * (n - 2), which is the general formula, for the answer plug 17 into n so 3 * (17 - 2)